LaTuyia:

A company sells boxes of bracelets (b) for $30 and boxes of necklaces (n) for $50. They make batches of bracelets that fill 5 boxes and batches of necklaces that fill 7 boxes. The company only has 48 boxes. Which system of equations helps the company plan to make $320? a b = 48 − n 5b + 7n = 48 b b = 48 − n 30b + 50n = 320 c 5b = 48 − 7n 30b + 50n = 320 d 5b = 48 − 7n 30b − 50n = 320

snowflake0531:

The variable b is boxes of bracelets ($30) The variable n is boxes of necklaces ($50) the amount of money amassed needs to equal 320 dollars 30b + 50n = 320 hopefully you understand how I got that They make batches of bracelets that fill 5 boxes and batches of necklaces that fill 7 boxes. The company only has 48 boxes. 5b + 7n = 48 -> 5b = 48 - 7n hopefully you understand how I got that too if you don't, please reply!